Solve the B.V.P.


𝑒𝑑 = 𝑒π‘₯π‘₯ + 𝑒; { 𝑒(π‘₯, 0) = cos π‘₯ 0 < π‘₯ < 1 𝑒(0,𝑑) = 0, 𝑒(1,𝑑) = 0Β 


1
Expert's answer
2021-11-01T17:36:46-0400

We use method of separating variables.

u(x.t)=X(x)"\cdot" T(t);

"T'(t)\cdot X(x)=T(t)\cdot X''(x)+T(t)\cdot X(x)"

Divide both parts by "T(t)\cdot X(x)"

"\frac{T'(t)}{T(t)}-1=\frac{X''(x)}{X(x)}"

Left part depend on t only but right part on x, it possible if both parts are equal to some constant "\lambda"

"\frac{T'(t)}{T(t)}-1=\frac{X''(x)}{X(x)}=\lambda"

"\begin{cases}
 X''(x)-\lambda\cdot X(x)=0; \\
 X(0)=0,\space X(1)=0.
\end{cases}"

This is standard bvp which has solution if "\lambda<0" and

"X(x)=C_1\cdot sin(\sqrt{-\lambda}\cdot x)+C_2\cdot cos(\sqrt{-\lambda}\cdot x)" -general solution.

Using boundary conditions we have:

"X(0)=C_2\cdot cos(0)=C_2=0\rarr X(x)=C_1\cdot sin(\sqrt{-\lambda}\cdot x)"

"X(1)=C_1\cdot sin(\sqrt{-\lambda})=0;
\sqrt{-\lambda}=\pi\cdot n,n=1,2,..."

"\lambda=-(\pi\cdot n)^2;"

Now we have set of solutions for n=1,2,...

"X_n(x)=C_n\cdot sin(\pi\cdot n\cdot x),n=1,2,..."

Further we shoulf find "T_n(t)" from an equation:

"\frac{T'(t)}{T(t)}-1=-(\pi\cdot n)^2;"

"\left(ln(T(t)) \right)'=1-(\pi\cdot n)^2"

"ln(T(t))=\left( 1-(\pi\cdot n)^2 \right)\cdot t+C"

The const C may be omitted because we have const Cn in Xn

Thus "T_n(t)=e^{\left( 1-(\pi\cdot n)^2 \right)\cdot t}=e^t\cdot e^{-(\pi\cdot n)^2\cdot t}"

For u(x,t) we use series

"u(x,t)=\sum_{n=1}^{\infty} C_nT_n(t)X_n(x)=\\e^{t}\cdot \sum_{n=1}^{\infty}C_ne^{-(\pi\cdot n)^2\cdot t} sin(\pi\cdot n\cdot x)"

Nou we take in account initial condition u(0,x)=cos(x) and have

"e^{0}\cdot \sum_{n=1}^{\infty}C_ne^{-(\pi\cdot n)^2\cdot 0} sin(\pi\cdot n\cdot x)=\\ \sum_{n=1}^{\infty}C_n sin(\pi\cdot n\cdot x)=cos(x)"

This is Furier series problem which has solution

"C_n=2\pi n\frac{cos(\pi n)\cdot cos(1)-1}{\pi^2n^2}"

Thus

u(x,t)= "2\pi e^{t}\cdot \sum_{n=1}^{\infty}\frac{cos(\pi n)\cdot cos(1)-1}{\pi^2n^2}e^{-(\pi\cdot n)^2\cdot t} sin(\pi\cdot n\cdot x)"